# Multi-scale Latent Point Consistency Models for 3D Shape Generation

### v2

@@ III-C Architecture Design
We now introduce the specific architecture of our noise prediction function $\epsilon_{\theta}$. Our backbone network is based on Point-Voxel CNNs (PVCNNs) [68]. To integrate multi-scale latent point feature, we propose the multi-scale latent integration (MLI) module before each SA and FP module of PVCNNs. We also adopt a 3D spatial attention mechanism in each point-voxel convolution (PVC) module, allowing the model to selectively attend to informative areas, as shown in Figure 2.
@@
where Norm is the layer normalization [69] and Linear denotes a linear layer. MLP consists of two linear layers. $\odot$ denotes element-wise product and $\circ$ means function composition. $\cdot\|\cdot$] means concatenation of vectors. Pos()($\cdot$) is the positional embedding function. To enable scale-time awareness in the model, we further modulate the feature with positional embedding of scale and time. The feature is fed to the following SA and FP modules to predict the noise.
3D Spatial Attention Mechanism. To help latent point representations better capture 3D spatial information , we design a 3D spatial attention mechanism. Our intuition is that closer points in 3D space would have a stronger correlation, thus higher attention values. We introduce a pairwise bias term that depends on the 3D distances between points. Specifically, in PVC, following [68], we first use PointNet++ [20] to extract the local features $F_{p}$ from the point cloud. We then convert the point cloud data into voxels, enabling the application of 3D convolution operations to obtain voxel features $F_{v}$. These are then fused to get the combined feature $F_{a}=\text{Add}(F_{p},\text{MLP}(F_{v}))$. For $F_{a}$, we compute the pairwise spatial distance matrix $B\in\mathbb{R}^{n^{i}\times n^{i}}$ between different point pairs at the current scale, where $n^{i}$ is the number of points at the current scale $i$, and use it as a bias in computing the attention,
@@
where $Q$, $K$, and $V$ are query, key, and value respectively in the standard attention module. We add this module to each PVC layer across all scales to make the model’s attention conform to the 3D distances between points. Additionally, following [17], we add 3D positional embeddings to adaptively aggregate feature from voxelized point clouds. In each PVC layer, given voxelized input $V\in\mathbb{R}^{v\times v\times v\times c}$, where $c$ is the embedding dimension and $v^{3}$ denotes the number of voxelized tokens, we add sine-cosine-based 3D positional embeddings to all input tokens.
@@ III-D Multi-Scale Latent Point Consistency Models
Since the diffusion models are notoriously slow in sampling, we explore the consistency models (CMs) [33] in the latent diffusion framework to accelerate the generation of 3D shapes. At the core of CMs, a consistency map is introduced to map any noisy data point on the trajectory of probability-flow ordinary differential equations (PF-ODEs) to the starting point, i.e., clean data. There are two ways to train such a consistency map [33], i.e., consistency distillation and consistency training. The former requires a pretrained teacher model whereas the latter trains the consistency map from scratch. We explored both options and found consistency distillation works well, whereas consistency training is sensitive to hyperparameters.
Specifically, we first rewrite the reverse process in the continuous-time setting following [36, 32],
@@
where $\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})$ is the noise prediction model and $\alpha_{t},\sigma_{t}$ are determined by the noise schedule as aforementioned. Samples can be drawn by solving the PF-ODE from $T$ to 0. To achieve consistency distillation, we intrsoduce a consistency map $f_{\hat{\theta}}:(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})\mapsto Z_{0}^{0}$ to directly predict the solution of PF-ODE in Equation 11. For $t=0$, we parameterize $f_{\hat{\theta}}$ using the noise prediction model $\hat{\epsilon}_{\theta}$:
@@
We assume that an effective ODE solver can be used to approximate the integral on the right side of Equation 11 from any starting time to the ending time. Note that we only use the solver during training and not during sampling. We aim to predict the solution of the PF-ODE by minimizing the consistency distillation loss [33]:
@@ IV Experiments
In this section, we train the multi-scale latent diffusion model on the Shapenet dataset and obtain MLPCM using the latent consistent distillation. We first introduce the dataset and evaluation metrics. Then we evaluate the performance of MLPCM on the Single-class 3D point cloud generation task. Next, we compare the results of MLPCM with a large number of baselines on multi-class unconditional 3D shape generation in terms of the performance, sampling time, and model parameters. Finally, we give a detailed ablation study on the effectiveness of invidual contributions.
Figure 3: Unconditional shape generation with 2,048 points for the classes airplane, car, and chair using class-specific models, trained on the ShapeNet dataset from PointFlow with global normalization. Our samples better represent component relationships, such as the seat-leg connection and the wing-fuselage connection, while competing approaches exhibit smaller gaps, fused components, or asymmetries between mirrored components.
Figure 4: Qualitative visualizations of high-fidelity and diverse 3D point cloud generation. Generated shapes preserve physically plausible structure and global topology while varying fine details, demonstrating diversity without sacrificing symmetry.
@@
Figure 6: Voxel-conditioned synthesis for interactive generation. Given a coarse voxel input, our model produces diverse 3D shapes that adhere to the occupied regions while refining geometry into clean surfaces and coherent parts.
@@ IV-A Datasets
For the unconditional 3D point cloud generation task, we utilize the ShapeNet dataset [70] as pre-processed and provided by PointFlow [2], ensuring consistency with prior studies. To facilitate fair comparisons with existing baseline methods [2, 71, 3, 17, 1], we focus on three commonly used categories: airplane, chair, and car. Each shape in the dataset contains 15,000 points, and during training, we randomly sample 2,048 points per shape in each iteration. The training dataset includes 2,832 shapes for airplanes, 4,612 for chairs, and 2,458 for cars.
We evaluate sample quality using a standard reference set, which consists of 405, 662, and 352 shapes for airplanes, chairs, and cars, respectively. Consistent with the normalization approach in PointFlow [2] and LION [3], the data is normalized globally across all shapes. Specifically, we calculate the mean values for each axis and a single standard deviation across all axes based on the entire training set. This ensures that the input data is standardized for training and comparison purposes.
While reproducing baseline methods on the ShapeNet dataset provided by PointFlow [2], we observed that certain approaches [3, 72, 73, 65, 74] utilize per-shape normalization. This involves calculating the mean for each axis individually for each shape and determining the scale based on the maximum extent across all axes for that shape. Consequently, the point coordinates are constrained within the range of [-1, 1]. For comparisons with these methods, we adhere to the same normalization strategy during both training and evaluation of our model. It is worth noting that these varying normalization schemes lead to fundamentally different generative modeling tasks. Therefore, ensuring clarity about these setups is crucial for fair and meaningful comparisons.
@@ IV-B Implementation Details
Training Details Our teacher model is implemented in PyTorch [75]. The VAE is trained for around 8000 epochs with an initial learning rate of 0.001, and we use Adam [76] to optimize the entire model. The batch size is set to 128, the latent point dimension is 4, and the skip weight is set to 0.01. The diffusion prior is trained for approximately 24,000 epochs with an initial learning rate of 0.0002, using 1000 time steps for the diffusion process, with a linear schedule for $\beta_{t}$. We train MLPCM for 100100K iterations with a batch size of $128$, a learning rate of $2\times 10^{-4}$, and an EMA rate $\mu=0.99995$. Both the VAE and the diffusion prior are trained on 8 NVIDIA A40 GPUs.
For the mixed denoising score network parametrization used in our teacher model, we follow the approach of [77]. Specifically, the denoising model is parameterized as an analytically optimal denoising network for a normal data distribution, augmented with a neural network-based correction. This design is particularly advantageous when the modeled distribution is close to normal, as is the case here. In the first training stage, all latent encodings are regularized to approximate a standard normal distribution due to the Kullback-Leibler regularization in the VAE objective. Our implementation of the mixed denoising score network strictly adheres to the method described by [77], and we direct readers to their work for additional details.
For the latent consistency model ,to align with the discrete-time schedule in our multi-scale latent diffusion model, which synthesizes shapes over 1000 DDPM steps, we employ DDIM [31] as the PF-ODE solver.
Evaluation Metrics Following recent works [2, 71, 3], we use 1-NNA, with both Chamfer distance (CD) and earth mover distance (EMD), as our main metric. It quantifies the similarity between the distribution of generated shapes and that of shapes from the validation set and measures both quality and diversity [2].
@@ IV-C Single-Class 3D Point Cloud Generation
Implementation Our implementation is based on the PyTorch. The input point cloud size is $2048\times 3$. Both our VAE encoder and decoder are built upon PVCNN [68], consisting of 4 layers of SA modules and FP modules. The voxel grid sizes of PVC at different scales are $32$, $16$, 88, and $8$, respectively. We adopt the Farthest Point Sampling (FPS) algorithm for sampling in the Grouper block, with the sampled center points being $1024$, $256$, $64$, and $16$. KNN (K-Nearest Neighbors) is used to aggregate local neighborhood features, with $32$ neighbors for each point in the Grouper. We apply a dropout rate of $0.1$ to all dropout layers in the VAE. Following [3], we initialize our VAE model so that it acts as an identity mapping between the input, latent space, and reconstruction points at the start of training. We achieve this by reducing the variance of the encoder and correspondingly weighting the skip connections. Our latent prior also consists of 4 layers of SA modules and FP modules. The time embedding and multi-scale latent representations are concatenated with the point features at the input of each corresponding SA and FP layer. We adopt the hybrid denoising score network parameterization following [3].
@@
TABLE I: Generation metrics (1-NNA↓) on airplane, chair, car categories from ShapeNet dataset from PointFlow [2]. Training and test data normalized globally into [-1, 1]. TM stands for Teacher Model and LCM stands for Latent Consistency Models.
@@
r-GAN [58] 98.40 96.79 83.69 99.70 94.46 99.01
l-GAN (CD) [58] 87.30 93.95 68.58 83.84 66.49 88.78
l-GAN (EMD) [58] 89.49 76.91 71.90 64.65 71.16 66.19
PointFlow [2] 75.68 70.74 62.84 60.57 58.10 56.25
SoftFlow [15] 76.05 65.80 59.21 60.05 64.77 60.09
SetVAE [78] 76.54 67.65 58.84 60.57 59.94 59.94
DPF-Net [16] 75.18 65.55 62.00 58.53 62.35 54.48
DPM [1] 76.42 86.91 60.05 74.77 68.89 79.97
PVD [71] 73.82 64.81 56.26 53.32 54.55 53.83
LION [3] 67.41 61.23 53.70 52.34 53.41 51.14
MeshDiffusion [79] 66.44 76.26 53.69 57.63 81.43 87.84
DiT-3D [17] 69.42 65.08 55.59 54.91 53.87 53.02
@@
TABLE II: Generation results (1-NNA↓) on ShapeNet dataset from PointFlow [2]. All data normalized individually into [-1, 1].
@@
Tree-GAN [60] 97.53 99.88 88.37 96.37 89.77 94.89
ShapeGF [65] 81.23 80.86 58.01 61.25 61.79 57.24
SP-GAN [72] 94.69 93.95 72.58 83.69 87.36 85.94
PDGN [74] 94.94 91.73 71.83 79.00 89.35 87.22
GCA [73] 88.15 85.93 64.27 64.50 70.45 64.20
LION [3] 76.30 67.04 56.50 53.85 59.52 49.29
@@
Tree-GAN [60] 96.80 96.60
PointFlow [2] 63.25 66.05
ShapeGF [65] 55.65 59.00
SetVAE [78] 79.25 95.25
PDGN [74] 71.05 86.00
DPF-Net [16] 67.10 64.75
DPM [1] 62.30 86.50
PVD [71] 58.65 57.85
LION [3] 51.85 48.95
@@
LION (DDPM) [3] 1000 27.09 53.41 51.14
LION (DDIM) [3] 1000 27.09 54.85 53.26
LION (DDIM) [3] 100 3.07 56.04 54.97
LION (DDIM) [3] 10 0.47 90.38 95.4
@@
TABLE VI: Ablation on the number of extra dimension $D_{l}$ of the latent points, on the car category.
@@
TABLE VII: Ablation study on the backbone of the encoder and decoder of hierarchical VAE and diffusion prior on the car category for Teacher Model. For DGCNN, we try different numbers of top-k nearest neighbors(knn).

### v3

@@ III-C Architecture Design
We now introduce the specific architecture of our noise prediction function $\epsilon_{\theta}$. Our backbone network is based on Point-Voxel CNNs (PVCNNs) [liu2019point]. To integrate multi-scale latent point feature, we propose the multi-scale latent integration (MLI) module before each SA and FP module of PVCNNs. We also adopt a 3D spatial attention mechanism in each point-voxel convolution (PVC) module, allowing the model to selectively attend to informative areas, as shown in Figure 2.
@@
where Norm is the layer normalization [ba2016layer] and Linear denotes a linear layer. MLP consists of two linear layers. $\odot$ denotes element-wise product and $\circ$ means function composition. $\cdot\|\cdot$] means concatenation of vectors. Pos()($\cdot$) is the positional embedding function. To enable scale-time awareness in the model, we further modulate the feature with positional embedding of scale and time. The feature is fed to the following SA and FP modules to predict the noise.
3D Spatial Attention Mechanism. To help latent point representations better capture 3D spatial information , we design a 3D spatial attention mechanism. Our intuition is that closer points in 3D space would have a stronger correlation, thus higher attention values. We introduce a pairwise bias term that depends on the 3D distances between points. Specifically, in PVC, following [liu2019point], we first use PointNet++ [qi2017pointnet++] to extract the local features $F_{p}$ from the point cloud. We then convert the point cloud data into voxels, enabling the application of 3D convolution operations to obtain voxel features $F_{v}$. These are then fused to get the combined feature $F_{a}=\text{Add}(F_{p},\text{MLP}(F_{v}))$. For $F_{a}$, we compute the pairwise spatial distance matrix $B\in\mathbb{R}^{n^{i}\times n^{i}}$ between different point pairs at the current scale, where $n^{i}$ is the number of points at the current scale $i$, and use it as a bias in computing the attention,
@@
where $Q$, $K$, and $V$ are query, key, and value respectively in the standard attention module. We add this module to each PVC layer across all scales to make the model’s attention conform to the 3D distances between points. Additionally, following [mo2023dit], we add 3D positional embeddings to adaptively aggregate feature from voxelized point clouds. In each PVC layer, given voxelized input $V\in\mathbb{R}^{v\times v\times v\times c}$, where $c$ is the embedding dimension and $v^{3}$ denotes the number of voxelized tokens, we add sine-cosine-based 3D positional embeddings to all input tokens.
@@ III-D Multi-Scale Latent Point Consistency Models
Since the diffusion models are notoriously slow in sampling, we explore the consistency models (CMs) [song2023consistency] in the latent diffusion framework to accelerate the generation of 3D shapes. At the core of CMs, a consistency map is introduced to map any noisy data point on the trajectory of probability-flow ordinary differential equations (PF-ODEs) to the starting point, i.e., clean data. There are two ways to train such a consistency map [song2023consistency], i.e., consistency distillation and consistency training. The former requires a pretrained teacher model whereas the latter trains the consistency map from scratch. We explored both options and found consistency distillation works well, whereas consistency training is sensitive to hyperparameters.
Specifically, we first rewrite the reverse process in the continuous-time setting following [song2020score, lu2022dpm],
@@
where $\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})$ is the noise prediction model and $\alpha_{t},\sigma_{t}$ are determined by the noise schedule as aforementioned. Sampling corresponds to integrating the PF-ODE backward from $T$ to 0. Following [luo2023latent, song2023consistency], we instead learn a consistency mapping $f_{\hat{\theta}}:(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})\mapsto Z_{0}^{0}$ that directly outputs the PF-ODE endpoint $Z^{0}_{0}$, enabling consistency distillation. At $t=0$, we express $f_{\hat{\theta}}$ in terms of the noise predictor $\hat{\epsilon}_{\theta}$:
@@
We assume a reliable ODE solver can approximate the integral in Eq. 11 over any time interval. The solver is used only during training, not at inference. We then learn to match the PF-ODE solution by minimizing a consistency distillation objective [song2023consistency]:
@@ IV Experiments
In this section, we train the multi-scale latent diffusion model on the Shapenet dataset and obtain MLPCM using the latent consistent distillation. We begin by describing the dataset and evaluation metrics, and then evaluate MLPCM on the single-class 3D point cloud generation benchmark. Next, we compare the results of MLPCM with a large number of baselines on multi-class unconditional 3D shape generation in terms of the performance, sampling time, and model parameters. Finally, we give a detailed ablation study on the effectiveness of invidual contributions.
Figure 3: We train class-specific ShapeNet models (airplane, car, chair) and generate 2,048-point unconditional samples with PointFlow-style global normalization. Our results better preserve part coherence such as seat–leg and wing–fuselage continuity, while prior methods often show disconnections, fused parts, or symmetry artifacts.
Figure 4: Qualitative results show high-quality and diverse 3D assets. Samples maintain realistic structure and overall topology while exploring fine-grained detail changes, yielding diversity with symmetry largely intact.
@@
Figure 6: Interactive voxel-guided generation. From a coarse occupancy grid, our model samples multiple plausible shapes that respect filled cells while sharpening geometry into smooth surfaces and well-formed parts.
@@ IV-A Datasets
For unconditional 3D point cloud generation, we adopt the ShapeNet benchmark [chang2015shapenet] using the official PointFlow preprocessing pipeline [yang2019pointflow] to keep the data protocol identical to prior work. For a controlled and fair comparison against representative baselines [yang2019pointflow, zhou20213d, zeng2022lion, mo2023dit, luo2021diffusion], we restrict evaluation to the three standard categories most frequently reported in the literature: airplane, chair, and car.
Each ShapeNet instance is provided as a dense point set with 15,000 points. During optimization, we perform on-the-fly subsampling by uniformly drawing 2,048 points per shape at every iteration, which acts as lightweight data augmentation and matches the common experimental setup. The resulting training split contains 2,832 airplane shapes, 4,612 chair shapes, and 2,458 car shapes, ensuring that all methods are trained and evaluated under the same category scope and point budget.
During our re-implementation of several baselines on the PointFlow ShapeNet splits [yang2019pointflow], we found that a subset of prior methods [zeng2022lion, li2021sp, zhang2021learning, cai2020learning, hui2020progressive] adopts a per-shape normalization protocol rather than a global, dataset-level normalization. Concretely, for each individual shape, they compute axis-wise means to center the point cloud, and then estimate a shape-specific scale using the maximum spatial extent across the three axes. After this centering-and-scaling step, all coordinates are mapped to a bounded cube, typically within [1,1][-1,1].
To ensure apples-to-apples comparisons with these methods, we follow the same per-shape normalization procedure in both training and evaluation when reporting results against that group. This detail is non-trivial as different normalization strategies effectively define different learning problems. Global normalization preserves category-level variations in absolute scale and aspect ratios, whereas per-shape normalization removes such information and emphasizes normalized geometry only. As a result, the results across the two protocols are not directly comparable unless the normalization setup is explicitly aligned. We therefore report and interpret numbers only within the same normalization regime to maintain fairness and avoid misleading conclusions.
@@ IV-B Implementation Details
Training Details We implement the teacher model in PyTorch [paszke2019pytorch] and optimize all components with Adam [kingma2014adam]. The VAE is trained for roughly 8,000 epochs, starting from a learning rate of $1\times 10^{-3}$ with a batch size of 128. We set the latent point dimensionality to 4 and use a skip weight of 0.01 to balance the reconstruction pathway during training. For the diffusion prior, we train for about 24,000 epochs with an initial learning rate of $2\times 10^{-4}$. The diffusion process uses 1,000 time steps with a linear schedule. We then train MLPCM for 100K iterations, again with batch size 128, a learning rate of $2\times 10^{-4}$, and an exponential moving average (EMA) decay of $\mu=0.99995$ to stabilize optimization and improve sampling quality. Both the VAE and diffusion prior are trained on 8 NVIDIA A40 GPUs. Finally, our mixed denoising score network is implemented to exactly match the formulations in [zeng2022lion, vahdat2021score].
Evaluation Metrics Consistent with prior studies [yang2019pointflow, zhou20213d, zeng2022lion], we evaluate using 1-NNA under both Chamfer Distance (CD) and Earth Mover’s Distance (EMD). This statistic compares the generated set against the validation distribution, jointly reflecting sample fidelity and diversity [yang2019pointflow].
@@ IV-C Single-Class 3D Point Cloud Generation
Implementation Our implementation is based on the PyTorch. The input point cloud size is $2048\times 3$. Both our VAE encoder and decoder are built upon PVCNN [liu2019point], consisting of 4 layers of SA modules and FP modules. The voxel grid sizes of PVC at different scales are $32$, $16$, 88, and $8$, respectively. We adopt the Farthest Point Sampling (FPS) algorithm for sampling in the Grouper block, with the sampled center points being $1024$, $256$, $64$, and $16$. KNN (K-Nearest Neighbors) is used to aggregate local neighborhood features, with $32$ neighbors for each point in the Grouper. We apply a dropout rate of $0.1$ to all dropout layers in the VAE.
@@
TABLE I: We evaluate generation on the airplane, chair, and car categories of the PointFlow ShapeNet split [yang2019pointflow] using 1-NNA (lower is better). We apply global normalization to map the train/test sets into [1,1][-1,1]. TM denotes the Teacher Model, and LCM denotes Latent Consistency Models.
@@
r-GAN [achlioptas2018learning] 98.40 96.79 83.69 99.70 94.46 99.01
l-GAN (CD) [achlioptas2018learning] 87.30 93.95 68.58 83.84 66.49 88.78
l-GAN (EMD) [achlioptas2018learning] 89.49 76.91 71.90 64.65 71.16 66.19
PointFlow [yang2019pointflow] 75.68 70.74 62.84 60.57 58.10 56.25
SoftFlow [kim2020softflow] 76.05 65.80 59.21 60.05 64.77 60.09
SetVAE [kim2021setvae] 76.54 67.65 58.84 60.57 59.94 59.94
DPF-Net [klokov2020discrete] 75.18 65.55 62.00 58.53 62.35 54.48
DPM [luo2021diffusion] 76.42 86.91 60.05 74.77 68.89 79.97
PVD [zhou20213d] 73.82 64.81 56.26 53.32 54.55 53.83
LION [zeng2022lion] 67.41 61.23 53.70 52.34 53.41 51.14
MeshDiffusion [liu2023meshdiffusion] 66.44 76.26 53.69 57.63 81.43 87.84
DiT-3D [mo2023dit] 69.42 65.08 55.59 54.91 53.87 53.02
@@
TABLE II: We report 1-NNA (lower is better) on the PointFlow ShapeNet benchmark [yang2019pointflow], using per-shape normalization to map each example into [1,1][-1,1].
@@
Tree-GAN [liu2018treegan] 97.53 99.88 88.37 96.37 89.77 94.89
ShapeGF [cai2020learning] 81.23 80.86 58.01 61.25 61.79 57.24
SP-GAN [li2021sp] 94.69 93.95 72.58 83.69 87.36 85.94
PDGN [hui2020progressive] 94.94 91.73 71.83 79.00 89.35 87.22
GCA [zhang2021learning] 88.15 85.93 64.27 64.50 70.45 64.20
LION [zeng2022lion] 76.30 67.04 56.50 53.85 59.52 49.29
@@
Tree-GAN [liu2018treegan] 96.80 96.60
PointFlow [yang2019pointflow] 63.25 66.05
ShapeGF [cai2020learning] 55.65 59.00
SetVAE [kim2021setvae] 79.25 95.25
PDGN [hui2020progressive] 71.05 86.00
DPF-Net [klokov2020discrete] 67.10 64.75
DPM [luo2021diffusion] 62.30 86.50
PVD [zhou20213d] 58.65 57.85
LION [zeng2022lion] 51.85 48.95
@@
LION (DDPM) [zeng2022lion] 1000 27.09 53.41 51.14
LION (DDIM) [zeng2022lion] 1000 27.09 54.85 53.26
LION (DDIM) [zeng2022lion] 100 3.07 56.04 54.97
LION (DDIM) [zeng2022lion] 10 0.47 90.38 95.4
@@
TABLE VI: We ablate the added latent-point feature dimension $D_{l}$ on the car category.
@@
TABLE VII: Ablation results on the teacher model on the car category evaluate the backbone choices in the hierarchical VAE and the diffusion prior. For DGCNN, we additionally test different $k$ values in the $k$-NN graph construction.